How do you solve this system of equations:
2x + 3y = 9
x - 2y = 1



Answer by rfer(16322)

2x+3y=9
x-2y=1
x=2y+1
2(2y+1)+3y=9
4y+2+3y=9
7y=9-2
7y=7
y=1
----------
2x+3*1=9
2x+3=9
2x=9-3
x=6/2
x=3
-----
(3,1)
------------
x-2y=1
3-2*1=1
1=1